OpenStudy (anonymous):

what is the derivative of g(x)=(1-2x)^2(x-3)

OpenStudy (anonymous):

looks like a product rule problem to me. plus some chain rule

OpenStudy (anonymous):

\[g(x)=(1-2x)^2(x-3)\] Use the product rule: \[g'(x)=(1-2x)^2\frac{d(x-3)}{dx}+\frac{d((1-2x)^2)}{dx}(x-3)\] Take the derivative where it says to, but don't forget about the chain rule: \[g'(x)=(1-2x)^2(1)+2*2(1-2x)(x-3)\] Clean it up a little... \[g'(x)=(1-2x)^2+4(1-2x)(x-3)\]

OpenStudy (anonymous):

Oh, shoot, I missed a negative sign when I did the chain rule. That should be a -2*2(...

OpenStudy (anonymous):

\[g'(x)=(1-2x)^2-4(1-2x)(x-3)\]

OpenStudy (anonymous):

why is it changing to - 4 i got \[+2(1-2x)^{-1}\],, is it wrong???

OpenStudy (anonymous):

Did that 2 come from the exponent or from the chain rule?

OpenStudy (anonymous):

exponent

OpenStudy (anonymous):

Ah, okay. So you need to use the chain rule: \[\frac{d((1-2x)^2)}{dx}=2(1-2x)^1\frac{d(1-2x)}{dx}\] \[=2(1-2x)^1(-2)=-4(1-2x)\]